Tatii:

What is 2x -12/4 =4??

mdobbs6856:

simplify 12/4 to 3 2x-3=4 add 3 to both sides 2x=7 Divide both sides by 2 7/2=x

mdobbs6856:

x=7/2

steevh:

@mdobbs6856 wrote:
x=7/2
yes !

Tatii:

I wish you could dum b it down for me bc i'm an idiot, but THANK YOU!!

mdobbs6856:

Well overall you're Welcome

Tatii:

:)

steevh:

she loves math ;)

Tatii:

0-o

Tatii:

not the wink-